OpenStudy (anonymous):

What is the area of the larger circle? (Round to the nearest tenth of a foot.)

OpenStudy (anonymous):



OpenStudy (anonymous):

truth be told I have no idea how to approach this... this somehow does not look like perfect circles... but if they were, is 3ft the radius of the small circle? and is 2ft the distance from the edge of the small circle to the big one?

OpenStudy (anonymous):

@hartnn need your help please

hartnn (hartnn):

is 3 ft radius or diameter of inner circle ??

OpenStudy (anonymous):

yes the radius is 3

hartnn (hartnn):

and is the distance between circles = 2 ft ???

OpenStudy (anonymous):

yes

hartnn (hartnn):

ok,then the radius of larger cirlce will be 3+2=5ft now the area of circle is given by \(A=pi \: r^2\) substiture r=5 and calculate the area.

OpenStudy (anonymous):

5 times 5 = 25

hartnn (hartnn):

and \(25 \pi =?\)

OpenStudy (anonymous):

i got 78.5 as my answer

hartnn (hartnn):

thats correct write it in ft.^2

OpenStudy (anonymous):

thankyou would you mind helping me with the next question?

hartnn (hartnn):

welcome:) ask.

OpenStudy (anonymous):

can you please help me with the next one